### v2

# Self-Generated Critiques Boost Reward Modeling for Language Models

Abstract
Reward modeling is crucial for aligning large language models (LLMs) with human preferences, especially in reinforcement learning from human feedback (RLHF). However, current reward models mainly produce unexplainable scalar scores and struggle to incorporate critiques in a natural language format. We hypothesize that generating both critiques and scalar rewards would improve reward models’ capability on preference ranking. Motivated by this, we propose Critic-RM, a framework that utilizes self-generated, high-quality critiques to train reward models for scalar reward-based preference prediction, with explicit rationales serving as supporting evidence. Critic-RM employs a two-stage process: generating and filtering high-quality critiques, followed by joint fine-tuning on reward prediction and critique generation objectives. Experiments on preference ranking benchmarks including RewardBench and CrossEval show that Critic-RM improves reward modeling accuracy by 3.7%–7.3% compared to standard reward models and LLM judges, demonstrating strong performance and data efficiency. Additional studies further validate the effectiveness of the generated critiques in rectifying flawed reasoning steps with the gain of 2.5%-3.2% on improving reasoning accuracy.
[…]
Yue Yu (), Rui Hou ()
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has been widely adopted to align large language models (LLMs) with human preferences (Ouyang et al., 2022; Touvron et al., 2023; Dubey et al., 2024; Reid et al., 2024). Central to the RLHF process is the reward model (RM), which is trained to assign scores that quantify how well the model’s outputs align with human judgments. The reward model defines optimization direction during training (e.g., reward signal in PPO), encouraging a policy LLM to generate more helpful, honest, and harmless responses ultimately enhancing the model’s generation quality in real-world applications.
Standard reward models are typically trained using preference pairs and optimized with pairwise logistic loss (Bradley and Terry, 1952), producing a single scalar score for each response. However, outputting a scalar score not only is hard to interpret but also fails to fully leverage the inherent language modeling capability that LLMs obtain from pretraining and post-training (Zhang et al., 2024). Consequently, these reward models tend to be less data-efficient and prone to robustness issues, such as reward hacking (Skalse et al., 2022; Singhal et al., 2023; Chen et al., 2024). Such limitations hinder the quality of feedback signals in RLHF and lead to suboptimal policy updates. On the other hand, the LLM-as-a-judge paradigm offers an alternative, where the LLM first generates a critique and then optionally provides a discrete score as a quality proxy for a response (Zheng et al., 2023; Kim et al., 2024a; Zhong et al., 2024). Combining the strengths of both paradigms—integrating the interpretability and structured critique of LLM-as-the-judge with the scalar optimization framework of reward models—has the great potential to address the limitations of each method and yield more robust and effective reward signals.
Despite its great premise, incorporating critiques into reward modeling presents two major challenges. (1) Conflicting objectives: Critique generation requires language modeling, while reward models provide scalar outputs, complicating its integration into language modeling. (2) Evaluator limitations: Off-the-shelf LMs are often not good evaluators, while additional fine-tuning requires costly human-generated or annotated critiques. Recent work (Ye et al., 2024) directly incorporates critiques generated from off-the-shelf LLMs for reward modeling, while Ankner et al. (2024) and Zhang et al. (2024) design a joint training approach for learning to generate the critique as well as rewards simultaneously via knowledge distillation. These methods typically rely on a strong teacher LLM to generate high-quality critiques, which can be costly and inefficient to obtain at scale in practice. Moreover, they cannot be used to improve frontier models when a stronger teacher model does not exist.
We introduce Critic-RM, a new framework that enhances reward models using synthetic critiques, without relying on strong LLM teachers. Our approach draws inspiration from recent advances in self-improving language models (Yuan et al., 2024; Wu et al., 2024; Prasad et al., 2024), where models are iteratively refined using data generated by themselves. To apply a similar LLM self-improving paradigm in reward modeling, we hypothesize that it is crucial to inject LLM’s critique generation ability into this process. Specifically, Critic-RM leverages an instruction-finetuned LLM as the backbone, which generates multiple candidate critiques, each with a discrete score (as explained below, for filtering critiques; not our final reward) for individual responses. However, these critiques can vary in quality, and poor-quality critiques often result in flawed quality predictions. To tackle this issue, we first apply a consistency-guided filtering technique, retaining only critiques whose scores align with human-annotated preference labels [^1] This discrete score is only used for filtering critiques and being different from the final reward score of Critic-RM. Our Critic-RM eventually produces a continuous score, as explained in Section 3.3.. To further enhance the quality of these synthetic critiques, we additionally propose two strategies, summarization and ranking, to refine the critiques used in training the reward model.
Once critiques are generated for each response, the main challenge lies in designing an effective training strategy to combine critique modeling and scalar reward prediction objectives. While LLMs benefit from learning through diverse critiques for each response (Ho et al., 2023), reward modeling is prone to overfitting (Dubey et al., 2024; Zhu et al., 2024); such a contradiction makes it nontrivial to determine the optimal learning steps. To address this issue, we introduce a simple weighting balancing strategy, where the model initially focuses on critique modeling loss, then gradually transitions to predicting rewards based on both the response and the critique. This approach balances the two learning objectives, allowing the model to excel at both high-quality critique generation and accurate reward prediction.
To demonstrate the effectiveness of Critic-RM, we conduct extensive experiments on RewardBench and three out-of-distribution reward modeling tasks, showing that Critic-RM outperforms baselines in both in-domain and out-of-domain evaluations. Additionally, experiments on critique evaluation benchmarks highlight Critic-RM’s ability to generate valuable feedback for correcting LLMs’ flawed reasoning. Our analysis confirms that Critic-RM’s superior generalization stems from its ability to identify and leverage high-quality self-generated critiques. The major contributions of our work can be summarized as follows:
  • We propose Critic-RM, a framework to allow LLMs to take advantage of self-generated critiques for reward modeling. Critic-RM does not rely on additional supervision compared to standard reward models, while enjoying an improved generation quality as well as reward modeling accuracy.
  • We propose a self-refinement technique to automatically select high-quality critiques, and design a simple yet effective weight scheduling strategy to balance the learning objectives between critique generation and reward modeling. These techniques collaboratively equip the model with the dual capabilities of high-quality critique generation and accurate reward prediction.
  • We conduct experiments on three benchmarks covering over ten tasks, demonstrating the effectiveness of Critic-RM in precise reward modeling across diverse scenarios. Additional studies confirm the utility of Critic-RM-generated critiques in identifying and correcting mistakes made by LLMs.
Table 1: Comparison of our proposed method Critic-RM and closest baselines.
Baselines Input Format Output Format Critique Generation Require Training Additional Teacher Models
Standard RM (Bradley and Terry) Single Response Continuous Score
RLAIF (Lee et al.) Single Response Continuous Score
LLM-as-a-judge (Zheng et al.) Response Pairs Discrete Score
SynRM (Ye et al.) Single Response + Critique Continuous Score
CLoud (Ankner et al.) Single Response Critique + Continuous Score
GenRM (Zhang et al.) Single Response Critique + Reward Token
Critic-RM (Ours) Single Response Critique + Continuous Score
2 Related Work
Reward Models. Building an accurate and robust reward model is a critical step for RLHF pipelines. Earlier work trains reward models with the ranking loss between chosen and rejected responses with the Bradley-Terry model (Bradley and Terry, 1952; Stiennon et al., 2020; Ouyang et al., 2022; Dubey et al., 2024). To further improve upon this reward modeling pipeline, Wang et al. (2024e, d, a) design fine-grained attributes to predict rewards toward different aspects, Chen et al. (2024); Shen et al. (2024b); Liu et al. (2024); Coste et al. (2024) promote the robustness of reward modeling via improved training techniques or model ensembling, and Pace et al. (2024); Shen et al. (2024a) study how to create synthetic examples for reward models. More related to us, several very recent works (concurrent to us) also study generative reward modeling. Ye et al. (2024) directly augment the response with additional critiques from a teacher model for reward modeling without training the RM for critique generation, and Zhang et al. (2024); Ankner et al. (2024); Mahan et al. (2024) attempt to learn reward models with additional critique objectives, with similar focus of our study. However, these methods typically rely on high-quality critiques from stronger teacher models for training, which can be costly and inefficient to obtain in practice. They also don’t provide a solution to reward modeling based on frontier LLMs where a teacher model doesn’t exist. They also lack a unified approach to improve the quality of the critiques. Besides, Zhang et al. (2024) is specific to verifying math problem correctness and is hard to map to subjective domains where there are no ground-truth answers.
LLM-as-a-judge and Critique Models. Recently, large language models (LLMs) have been proposed as cost-effective alternatives to human evaluation, and act as proxies for assessing text quality. Such methods often first provide explanations for judgments of the response, then output a discrete score or preference label as the prediction (Zheng et al., 2023; Li et al., 2023; Yan et al., 2024; Xu et al., 2024). CriticGPT (McAleese et al., 2024) has also extended this line of work into coding tasks, where the LLM critic model is fine-tuned to pinpoint problems in code from real-world assistant tasks. However, using off-the-shelf LLMs for evaluation introduces the risk of bias (Bavaresco et al., 2024; Stureborg et al., 2024), and they can be easily misled (Zeng et al., 2024). To address these challenges, recent studies (Wang et al., 2024b; Kim et al., 2024a) have focused on collecting high-quality response pairs to train more accurate and reliable LLM-based evaluators.
Self-alignment Techniques. Aligning LLMs with human preferences often requires a massive amount of human annotations. To alleviate this reliance on human efforts, self-alignment leverages the model’s own capabilities to refine its responses and align them with desired behaviors. Saunders et al. (2022); Madaan et al. (2023) use LLM itself to refine the original response at the inference time. Li et al. (2024b) generate instruction prompts for web documents and subsequently select high-quality examples for instruction fine-tuning. Lee et al. (2024); Sun et al. (2024) leverage LLMs to create preference labels efficiently, Yuan et al. (2024) employ LLM itself to rank different responses to provide its own rewards during training, and Zelikman et al. (2022); Pang et al. (2024); Gulcehre et al. (2023) improve LLM reasoning abilities through self-generated reasoning steps. A recent study (Wang et al., 2024b) also employs self-improving techniques to train text evaluators, but it focuses on pairwise evaluation and generating synthetic preference pairs. In contrast, we combine self-generated critiques with human-annotated preference pairs to enhance reward modeling performance.
3 Methodology
3.1 Preliminaries
Reward Modeling. Let $\mathcal{X}$ and $\mathcal{Y}$ denote the space of prompts and responses, respectively. In the RLHF pipeline, human feedback is typically collected in the form of pairwise preferences between two responses $\left(y^{+},y^{-}\right)\in\mathcal{Y}^{2}$ to a given prompt $x\in\mathcal{X}$. Then, the preference dataset can be written as $\mathcal{D}=\left\{(x_{i},y_{i}^{+},y_{i}^{-})\right\}_{i=1}^{|\mathcal{D}|}$, where the preference for $y^{+}$ over $y^{-}$ is denoted as $y^{+}\succ y^{-}$. To model the pairwise preferences, the learning objective is to maximize the probability with Bradley-Terry model (Bradley and Terry, 1952) as
$p\left(y^{+}\succ y^{-}\mid x\right)=\frac{\exp\left(r\left(x,y^{+}\right)\right)}{\exp\left(r\left(x,y^{+}\right)\right)+\exp\left(r\left(x,y^{-}\right)\right)}.$ (1)
In practice, the reward model $r_{\psi}$ is trained to minimize the following empirical negative log-likelihood loss (Stiennon et al., 2020; Ouyang et al., 2022; Dubey et al., 2024):
$\ell_{\text{rm}}(\psi)=-\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim\mathcal{D}}\log\left(\sigma\left(r_{\psi}\left(x,y^{+}\right)-r_{\psi}\left(x,y^{-}\right)\right)\right.$ (2)
where $\sigma$ denotes the sigmoid function.
Problem Setup. In this work, we investigate the usage of off-the-shelf instruction-finetuned LLM $\mathcal{M}_{\theta}$ as the backbone for both the critique generation model and reward model. Specifically, we denote the critic generation model as $g_{\phi}=h_{\text{g}}\circ\mathcal{M}_{\theta}$ and the reward model as $r_{\psi}=h_{\text{r}}\circ\mathcal{M}_{\theta}$, where $h_{\text{g}}$ and $h_{\text{r}}$ stand for the language modeling head (inherited from the original $\mathcal{M}_{\theta}$) and reward modeling head (randomly initialized).
Figure 1: An overview of Critic-RM. For each preference pair in the training set, we begin by prompting the LLM to generate candidate critiques along with discrete scores. Next, instance-level critique filtering is applied to minimize the impact of examples that conflict with preference labels. Finally, quality-aware critique refinement is performed to produce critiques that enhance reward model training.
Overview of Critic-RM. The framework of Critic-RM is shown in Figure 1. Critic-RM first generates candidate critiques for each prompt-response pair. Then, a filtering step is conducted to reduce the effect of potentially noisy rationales leading to incorrect predictions, allowing us to augment the preference pairs with additional critiques with the goal of improving the precision of reward modeling. Finally, we implement a joint training scheme to teach the model both high-quality critique generation and accurate reward modeling. The following sections will provide more details about each step.
3.2 Critique-augmented Reward Model Training
To integrate the critiques into the reward modeling step, we view critiques as latent variables, which serve as an intermediate variable between the response and the final reward. Specifically, we denote $z^{+},z^{-}$ as critiques for chosen and rejected responses $y^{+},y^{-}$ with prompt $x$, respectively. Then, the overall learning objective $p\left(y^{+}\succ y^{-}\mid x\right)$ can be recast as
$p(y^{+}\succ y^{-}\mid x)$ $=\sum_{z^{+},z^{-}}p(y^{+}\succ y^{-},z^{+},z^{-}\mid x)$ (3)
$=\sum_{z^{+},z^{-}}p(y^{+}\succ y^{-}\mid z^{+},z^{-},x)\cdot p^{*}(z^{+}\mid y^{+},x)\cdot p^{*}(z^{-}\mid y^{-},x).$
Since $p^{*}(\cdot\mid y,x)$ stands for the oracle distribution for critiques and is often not intractable, we aim to leverage the critic generation model $g_{\phi}$ to generate the approximate distribution $q_{\phi}$ by applying the Jensen’s Inequality as
$\log p\left(y^{+}\succ y^{-}\mid x\right)$ $=\log\mathbb{E}_{q_{\phi}\left(z^{+}\mid y^{+},x\right),q_{\phi}\left(z^{-}\mid y^{-},x\right)}\left[\frac{p\left(y^{+}\succ y^{-},z^{+},z^{-}\mid x\right)}{q_{\phi}\left(z^{+}\mid y^{+},x\right)q_{\phi}\left(z^{-}\mid y^{-},x\right)}\right]$ (4)
$\geq\mathbb{E}_{q_{\phi}\left(z^{+}\mid y^{+},x\right),q_{\phi}\left(z^{-}\mid y^{-},x\right)}\left[\log\frac{p\left(y^{+}\succ y^{-},z^{+},z^{-}\mid x\right)}{q_{\phi}\left(z^{+}\mid y^{+},x\right)q_{\phi}\left(z^{-}\mid y^{-},x\right)}\right]$
Then, instead of directly optimizing the negative log-likelihood, the training objective can be expressed as
$\mathcal{L}$ $=\mathbb{E}_{q_{\phi}\left(z^{+}\mid y^{+},x\right),q_{\phi}\left(z^{-}\mid y^{-},x\right)}\left[-\log\frac{p\left(y^{+}\succ y^{-},z^{+},z^{-}\mid x\right)}{q_{\phi}\left(z^{+}\mid y^{+},x\right)q_{\phi}\left(z^{-}\mid y^{-},x\right)}\right]$ (5)
$=\underbrace{\mathbb{E}_{{q_{\phi}}\left(z^{+}\mid y^{+},x\right),q_{\phi}\left(z^{-}\mid y^{-},x\right)}\left[-\log p\left(y^{+}\succ y^{-}\mid z^{+},z^{-},x\right)\right]}_{\text{Preference Modeling Loss with Critiques}}$
$+\underbrace{\mathcal{D}_{\text{KL}}\left((q_{\phi}(z^{+}\mid y^{+},x)\|p^{*}(z^{+}\mid y^{+},x)\right)+\mathcal{D}_{\text{KL}}\left((q_{\phi}(z^{-}\mid y^{-},x)\|p^{*}(z^{-}\mid y^{-},x)\right)}_{\text{Critique Generation Loss}}.$
What does the learning objective imply? Eq. 5 provides a way to decompose the reward model learning objective into two parts: (1) Preference Modeling Loss with Critiques $\ell_{\text{r}}$: the reward model $r_{\theta}$ learns to predict the reward for each response conditioned on critiques; (2) Critique Generation Loss $\ell_{\text{c}}$: the LLM generation $g_{\theta}$ is trained to generate critiques to approximate the oracle distribution $p^{*}(\cdot\mid y,x)$. We will discuss how to train the reward model $r_{\theta}$ and critique generation model $g_{\theta}$ in the following subsections.
3.2.1 Critique-augmented Reward Prediction
To enable the reward model $r_{\psi}$ to learn the preference with critiques (i.e., $\ell_{\text{r}}$) can be straightforward, as we only need to modify the input by augmenting response with critiques as
$\ell_{\text{r}}(x,y^{+},y^{-},z^{+},z^{-})=-\log p\left(y^{+}\succ y^{-},z^{+},z^{-}\mid x\right)=-\log p\left(r_{\psi}(x,[y^{+};z^{+}])>r_{\psi}(x,[y^{-};z^{-}])\right).$ (6)
In this way, for each prompt, the reward model will learn to generate the reward based on both responses and critiques. In practice, we put the critiques after the response and add a special token at the end of the critique for calculating the reward.
3.2.2 Critique Generation & Filtering
For critique generation loss, approximating $p^{*}(\cdot\mid y,x)$ can be nontrivial as the primary challenge lies in the lack of high-quality critique annotations. To ensure the quality of the critiques, our key hypothesis is that good critiques for responses should align well with human preference labels. With this in mind, we design a generate-then-filter framework to create high-quality supervision signals for critique model training.
Critique Generation. To generate critiques without relying on stronger LLMs, we first prompt the LLM $\mathcal{M}_{\theta}$ (with the same backbone as the reward model) and sample a set of $N$ candidate critiques for input prompt and responses $(x,y)$ by following the procedure of the LLM-as-a-judge pipeline as $(\widehat{z}_{i},s_{i})_{i=1}^{N}\sim g_{\phi}(x,y)$, where $\widehat{z}$ is the generated critique and $s$ is a discrete score ranging from 1 to 10, indicating the quality of the response.
Instance-level Critique Filtering. To reduce the potential noisy critiques and encourage the consistency between critiques and preference labels, we propose to first retain instances guided by the score generated by the judge in the previous score as $\mathcal{D}_{\text{sub}}=\{(x,y^{+},y^{-})\mid\bar{s}(x,y^{+})>\bar{s}(x,y^{-})\}$, where $\bar{s}(x,y^{+})=\sum_{i=1}^{N}s_{i}^{+}/{N}$ and $\bar{s}(x,y^{-})=\sum_{i=1}^{N}s_{i}^{-}/{N}$ stand for the average score for chosen and rejected responses, respectively. By applying this filtering process, we enhance the consistency of critiques with human preferences and minimize the impact of noisy instances.
Quality-aware Critique Refinement. The previous step mainly focuses on instance-level denoising, while for each (prompt, response) pair, the quality of different critiques also varies. To further improve the quality of critiques, we design a Meta-judge-based technique (Wu et al., 2024) to leverage LLM $\mathcal{M}_{\theta}$ again to further refine the critiques in $\mathcal{D}_{\text{sub}}$, with two possible variants:
  • Summarization-based Refinement: We adopt the LLM as a summarizer to write ‘meta-critiques’ given different critiques so that the LLM can potentially identify the most common, reasonable feedback while mitigating the impact of the potential incorrect feedback. The final critique can be written as $\mathcal{Z}_{\text{summ}}=\{z_{i}\}_{i=1}^{K}\sim g_{\phi}(x,y,\Pi_{j=1}^{N}{\widehat{z}_{j}})$, where $\Pi_{j=1}^{N}{\widehat{z}_{j}}$ is a permutation of $N$ initial critiques. By sampling over different permutations of critiques, we can generate more diverse critiques for model training.
  • Ranking-based Refinement: We use the LLM as a meta-judge to create evaluation scores for critiques. Specifically, for each critique $\widehat{z}_{i}$, we prompt the LLM to generate a discrete score from 1 to 10 as $m_{i}\sim g_{\phi}(x,y,\widehat{z}_{i})$, which serves as a proxy for critique quality estimation. Then, we only retain top-$K$ ranked critiques as $\mathcal{Z}_{\text{rank}}=\{z_{i}\}_{i=1}^{K}=\operatorname{Top-K}(\{\widehat{z}_{i}\}_{i=1}^{N})$. In this way, we can preserve the critiques with the highest quality identified by the model itself.
Final Loss for Critique Generation. From the previous step, we augment the training set $\mathcal{D}_{\text{sub}}$ with self-identified high-quality critiques, denoted as $\mathcal{D}_{\text{sub}}=\{(x,y^{+},y^{-},\mathcal{Z}^{+},\mathcal{Z}^{-})\}$. With the self-generated high-quality critiques $\mathcal{Z}$, we aim to use them to approximate the distribution of oracle distribution as $p^{*}(z\mid y,x)=\mathbb{I}(z\in\mathcal{Z})$. Directly using this distribution in backward KL loss in Eq. 5 may lead to policy and entropy collapses (Sessa et al., 2024; Agarwal et al., 2024). As a result, we use forward KL loss to approximate this learning objective. Then using the empirical distribution, the KL divergence becomes:
$\ell_{\text{c}}(\mathcal{Z};x,y)$ $=\mathcal{D}_{\mathrm{KL}}(p^{*}(z\mid y_{i},x_{i})\|q_{\phi}(z\mid y_{i},x_{i}))$ (7)
$=\mathbb{E}_{z\sim p^{*}(z\mid y_{i},x_{i})}\left[p^{*}(z\mid y_{i},x_{i})-\log q_{\phi}\left(z\mid y_{i},x_{i}\right)\right]$
$=-\frac{1}{K}\sum_{z\in\mathcal{Z}}\log q_{\phi}(z\mid y,x)+\text{const}.$
Then, the overall loss for critique generation can be written as $\ell_{\text{c}}(x,y^{+},y^{-},\mathcal{Z}^{+},\mathcal{Z}^{-})=\ell_{\text{c}}(\mathcal{Z}^{+};x,y^{+})+\ell_{\text{c}}(\mathcal{Z}^{-};x,y^{-}).$
3.2.3 Joint Learning of Critique Generation and Reward Modeling
To combine the reward modeling loss (Eq. 6) and critique generation loss together (Eq. 7), one challenge lies in the different learning objectives for these two terms: for critique generation, the model $g_{\phi}$ will benefit more from fine-tuning with diverse critiques from $\mathcal{Z}$. On the contrary, the reward model $r_{\psi}$ is often observed with overfitting issues when fine-tuning with more than one round. To resolve this issue, we design a dynamic weight schedule approach, where we add an additional weight $\lambda(t)$ on Equation 5, which is relevant to the training step $t$, for balancing between these two objectives as
$\mathcal{L}(\phi,\psi)=\mathbb{E}_{(x,y^{+},y^{-},\mathcal{Z}^{+},\mathcal{Z}^{-})\in\mathcal{D}_{\text{sub}}}\left[\lambda(t)\cdot\ell_{c}(\phi)+(1-\lambda(t))\cdot\ell_{\text{r}}(\psi)\right],$ (8)
where $\lambda(t)$ is defined as
$\lambda(t)=\begin{cases}1,&{0<t<(K-1)T}\\ 1-\beta\times\frac{t-(K-1)T}{T}.&{(K-1)T<t<KT}\end{cases}$ (9)
Here, $T$ represents the total number of training steps in one epoch. This approach allows the model to focus on critique generation during the initial phase of training and shifts to reward learning in the final round, mitigating the overfitting issue in the reward model.
3.3 Critic-RM Inference
Compared to standard reward model training, Critic-RM involves an additional step for each (prompt, response) pair during inference. Specifically, given the (prompt, response) pair $(x,y)$, the model will first generate a critique $z\sim q_{\phi}(x,y)$, then predict the reward for the response as $r=r_{\psi}(x,[y,z])$.
Inference-time Scaling. Following recent studies (Ankner et al., 2024; Zhang et al., 2024), we also conduct inference-time scaling (Wang et al., 2023) to improve performance. Specifically, we generate a set of $m$ critiques as $\mathcal{Z}=\{z_{i}\}_{i=1}^{m}\sim q_{\phi}(x,y)$ with non-zero temperatures, then predict the reward for the response as the average of reward over different critiques as $r=r_{\psi}(x,[y,z_{i}])/m$.
4 Experiments
4.1 Experiment Setup
4.1.1 Training Data
To ensure the representativeness of the preference pairs used in this study, we leverage both public and synthetic datasets for reward model training.
Public Preference Datasets: We choose a set of datasets for reward model training with human-generated preference labels mainly from public, open-sourced datasets (Ivison et al., 2024; Wang et al., 2024a). We include the following datasets:
  • General Chat Domain: We include datasets from ChatArena (Zheng et al., 2023) and AlpacaFarm-Human-Pref (Dubois et al., 2023).
  • Helpfulness Data: We leverage HelpSteer2 (Wang et al., 2024d) to create preference data.
  • Reasoning: We mainly use Evol-instruct (Xu et al., 2023) which contains preference pairs for complex instruction following, coding-related tasks.
  • Safety: We employ PKU-SafeRLHF (Dai et al., 2024), which includes safety-related prompts paired with both safe and unsafe responses to form preference pairs.
Synthetic Preference Datasets: To incorporate additional preference supervision from different domains, we further include synthetic data using Llama-3.1 models. [^2] These synthetic data are used for both Critic-RM and our direct baselines. Specifically, for the math domain, we consider questions in GSM8K (Cobbe et al., 2021) and the MATH dataset (Hendrycks et al., 2021). For each math question, we use Llama-3.1-8b-instruct, and Llama-3.1-70b-instruct to generate candidate solutions with the prompt “Given the following problem, reason step-by-step and give a final answer to the problem.”, and generate multiple candidate solutions for a given prompt. We use those responses that lead to correct solutions as the chosen response while considering those responses with incorrect solutions as the rejected response. In the safety domain, we generate synthetic prompts following the safety principles outlined in SafeRLHF (Dai et al., 2024) (e.g., Hate Speech, Offensive Language, Discrimination, Violence). To ensure balance, we also include scenarios where the model should not refuse to respond (e.g., Figurative Language, Safe Targets testing for ambiguous meanings) to avoid skewing the data toward over-conservatism.
4.1.2 Evaluation Benchmarks.
Evaluation Benchmarks for Reward Models. In our experiments, we mainly evaluate on RewardBench (Lambert et al., 2024), which contains a collection of prompt-chosen-rejected triplets across chat, reasoning, and safety domains, including 2985 examples in total. We use the standard evaluation protocol provided by the original authors. Beyond RewardBench, we also aim to test the out-of-distribution generalization ability of reward models. Specifically, we consider CrossEval (Zhong et al., 2024), a recently proposed benchmark to evaluate the LLM’s capability in real-world interactions [^3] The details for data processing is listed in Appendix 6.. Besides, we also consider two additional datasets, namely QA Feedback (Wu et al., 2023) and SHP (Ethayarajh et al., 2022), which focuses on evaluating the response for open-ended QA task as well as social platforms (i.e., Reddit). There are around 2000 examples of QA Feedback preference pairs. For SHP, we use the response with a higher average score/votes judged by human raters as the positive response, and we use the response with a lower score/votes as the negative one, and randomly subsample 3000 pairs for evaluation. For all tasks, we use accuracy as the main metric.
Evaluation Benchmarks for Critic Models. To demonstrate the effectiveness of our model in generating improved critiques, we employ CriticBench (Lin et al., 2024), a benchmark designed to evaluate LLMs’ ability to critique and improve their reasoning across various tasks. CriticBench covers five key reasoning domains: mathematical, commonsense, symbolic, coding, and algorithmic. It includes responses from 17 different LLMs, requiring the LLMs to provide critiques that assess the correctness of these LLMs’ responses. Specifically, it considers two dimensions for evaluation: (1) Critique Accuracy: where F1 Score is used to evaluate the correctness of critiques; (2) Correction Accuracy: where Accuracy is used to evaluate whether the model can generate correct answers based on critique feedback.
4.1.3 Baselines
We consider the following baselines from three different groups:
  • LLM-as-a-judge: With the prompt with a pair of responses used as the input, this line of models needs to generate a preference label. We consider Prometheus-v2 (Kim et al., 2024b), Llama-3.1-70B/405B (Dubey et al., 2024), GPT-4 (Achiam et al., 2023) and GPT-4o (Hurst et al., 2024), Gemini-1.5-pro (Reid et al., 2024) and recently proposed self-taught evaluator (Wang et al., 2024b) based on Llama-3-70B for comparison.
  • Standard Reward Models: This line of models only outputs a scalar score for each (prompt, response) pair. We compare with baselines including standard RM (Stiennon et al., 2020), Cohere-0514, SteerLM-RM (Wang et al., 2024e), Nemotron-RM (Adler et al., 2024).
  • Reward Model with Critiques: These studies are mostly relevant to us as they also leverage critiques to improve reward models. Specifically, we compare with SymRM (Ye et al., 2024) which directly augments responses with critiques for reward modeling, and CLoud (Ankner et al., 2024) which jointly learn to generate critiques and predict rewards.
It is worth noting that for most relevant baselines (e.g. RM, SynRM, CLoud), we reimplement those baselines with the same training data and backbone to ensure the comparison is fair and meaningful. We do not consider some reward model training techniques (Wang et al., 2024c, a) as they focus on designing better learning objectives for standard reward models, which are orthogonal to the focus of this study.
4.1.4 Implemenation Details
We use Llama3.1-70B-Instruct (Dubey et al., 2024) as the backbone in our main experiments. For critique generation, we set the temperature $\tau=0.9$ and sample $N=10$ candidate critiques for each response. For the critique filtering, we set $K=2$ to select top-2 responses. For model fine-tuning, we use the Adam optimizer (Kingma and Ba, 2014) with the learning rate 2e-6, weight decay 0.1 and dropout 0.1. We set the global batch size to 64, $\beta$ in Eq. 9 to 0.9 and train the model with $2$ epochs. We observe that there exist several examples in AlpacaEval and ChatArena that share similar prompts with the target evaluation tasks, and we remove all overlapping prompts to avoid data contamination (Oren et al., 2024). During inference, if inference-time scaling is adopted, we choose temperate $\tau=0.95$ to sample multiple critiques. The prompt format we use in experiments is exhibited in Appendix 7.
Table 2: Results of our proposed method and baselines on the RewardBench. : Results copied from either RewardBench Leaderboard or original papers. §: This version of the model is trained using SFT only.
Models Chat Chat_Hard Reasoning Safety Overall
LLM-as-a-judge (For Reference)
Prometheus-8*7b-v2 (Kim et al., 2024b) 93.0 47.1 77.4 80.5 74.5
Llama3.1-70B-Instruct (Dubey et al., 2024) 97.2 70.2 82.8 86.0 84.0
Llama3.1-405B-Instruct (Dubey et al., 2024) 97.2 74.6 77.6 87.1 84.1
GPT-4-0125 (Achiam et al., 2023) 95.3 74.3 87.6 86.9 86.0
GPT-4o-0806 (Hurst et al., 2024) 96.1 76.1 88.1 86.6 86.7
Gemini-1.5-pro-0514 (Reid et al., 2024) 92.3 80.6 92.0 87.9 88.2
Self-taught Evaluator§ (Wang et al., 2024b) (Iter 1) 98.3 69.0 82.6 85.7 83.9
Self-taught Evaluator§ (Wang et al., 2024b) (Iter 2) 97.5 75.4 81.7 89.5 86.0
Self-taught Evaluator§ (Wang et al., 2024b) 96.6 84.2 91.5 81.0 88.3
    w/ inference scaling, $m=32$ 96.9 84.0 91.5 82.5 88.7
Standard Reward Models
RM (Stiennon et al., 2020) 98.3 74.5 88.0 83.8 86.4
Cohere-0514 96.4 71.3 92.3 97.7 89.4
SteerLM-RM 70B (Wang et al., 2024e) 91.3 80.3 92.8 90.6 88.8
Nemotron-RM 340B (Adler et al., 2024) 95.8 87.1 91.5 93.6 92.0
(Concurrent Work) Reward Models with Critiques
SynRM (Ye et al., 2024) (Reported Best) 38.0 82.5 87.1 74.1 70.4
SynRM (Ye et al., 2024) (Ours) 97.5 76.8 88.5 86.3 87.3
CLoud (Ankner et al., 2024) (Reported) $\sim$97.0 $\sim$58.0 $\sim$92.0 $\sim$84.0 82.8
CLoud (Ankner et al., 2024) (Ours) 98.0 75.6 87.6 89.0 87.6
    w/ inference scaling, $m=32$ 98.0 75.2 89.3 91.5 88.5
Critic-RM-Summ 98.0 77.0 88.9 94.5 89.6
    w/ inference scaling, $m=32$ 97.5 77.0 91.6 95.9 90.5
Critic-RM-Rank 97.5 79.6 90.6 94.1 90.5
    w/ inference scaling, $m=32$ 97.2 80.0 91.6 95.1 91.0
4.2 Main Experiments: RewardBench
Table 2 presents results of Critic-RM and baselines. The findings are summarized as follows:
  • Incorporating Critiques Helps Reward Modeling in General. Critic-RM generally outperforms the baselines used in this study. Specifically, when trained with the same preference data, Critic-RM outperforms the standard Reward Model by 3.7%-4.7%. Critic-RM also outperform giant Llama-3.1-405b judge model by 6.2%-7.3%, respectively. These results justify the advantage of incorporating critiques into the reward model training step, which facilitates both high-quality critiques and precise rewards.
  • High-quality Critiques Matters. By comparing Critic-RM with baselines that also incorporate critiques into reward modeling, we observe that their performance gains over the standard RM are smaller than ours. We attribute this performance gap to the lack of post-processing methods for improving critique quality, which is key to achieving self-improvement in this challenging setting.
  • Inference-time Scaling Mainly Helps for Reasoning Tasks. We observe further performance improvements for both Critic-RM and the baselines when multiple critiques are generated during inference. Notably, these gains are most pronounced in reasoning-intensive tasks such as Math, Coding, and Safety, where the model must decide whether to reject a response. This suggests that, when computational resources are constrained, prioritizing reasoning-heavy tasks can lead to more significant performance improvements.
Table 3: Results of our proposed method and baselines on out-of-distribution reward modeling datasets.
Models CrossEval Other Datasets
English Reasoning Coding Tool C+R T+R T+C Avg. QA Feedback SHP
LLM-as-a-judge (For Reference)
Llama3.1-70B-Instruct (Dubey et al., 2024) 55.4 71.4 70.1 77.4 78.2 69.5 80.7 71.8 59.2 63.3
Llama3.1-405B-Instruct (Dubey et al., 2024) 64.4 71.9 77.5 80.2 78.2 75.6 78.9 75.2 60.7 62.9
Reward Models
RM (Stiennon et al., 2020) 59.3 72.7 70.8 75.2 68.3 72.0 72.4 70.1 58.3 65.1
CLoud (Ankner et al., 2024) 60.3 75.2 71.7 79.0 73.2 71.1 73.4 72.0 59.2 64.8
Our Model
Critic-RM-Summ 61.3 76.2 72.4 80.7 73.2 71.6 76.9 73.0 60.4 67.9
Critic-RM-Rank 64.0 74.3 73.3 80.7 79.3 72.0 79.3 74.7 60.2 66.2
4.3 Out-of-Distribution (OOD) Evaluation
As shown in Table 3, we evaluate the performance of our approach (Critic-RM) alongside relevant baseline models on three out-of-distribution (OOD) reward modeling datasets. Our results demonstrate that Critic-RM exhibits a strong performance across these datasets, surpassing standard reward modeling (RM) baselines by an average margin of 4%. Notably, the performance improvements of Critic-RM are more pronounced on more challenging benchmarks, such as tasks requiring cross-abilities, suggesting that the benefits of critiques are more significant in complex scenarios. Furthermore, we observe that the performance of Critic-RM is comparable to that of LLM-judge models with significantly more parameters. This highlights the efficiency and effectiveness of Critic-RM when being adapted to real scenarios.
Table 4: Results of our proposed method and baselines on CriticBench (Lin et al., 2024). : For these methods, we use the same Llama-3.1-8b-Instruct as the backbone model for answer correction.
Models Critique Accuracy (F1) Correction Accuracy (Acc.)
Algorithm Code Symbolic Commonsense Math Total Algorithm Code Symbolic Commonsense Math Total
Baselines
Auto-J 13B (Li et al., 2024a) 65.29
UltraCM 13B (Cui et al., 2024) 61.11
CLoud (Ankner et al., 2024) 57.22 82.87 80.56 70.18 90.35 81.91 84.75 74.56 95.35 50.22 68.48 69.56
GPT-3.5 (OpenAI, 2022) 46.15 73.13 64.49 50.22 62.01 61.11 58.16 61.85 71.83 44.11 41.95 51.24
GPT-4 (Achiam et al., 2023) 63.51 91.36 90.75 71.56 92.55 78.75 77.66 76.29 92.41 59.96 63.57 69.96
LLM-as-a-judge (For Reference)
Llama3.1-70B-Instruct (Dubey et al., 2024) 60.37 84.92 86.17 65.52 88.53 80.75 77.65 76.93 88.06 59.29 57.28 66.96
Llama3.1-405B-Instruct (Dubey et al., 2024) 86.96 88.96 90.70 72.59 93.84 86.96 86.52 81.42 90.86 63.76 63.36 72.02
Our Model
Critic-RM-Summ 89.79 89.36 88.36 75.26 96.09 88.25 90.55 81.89 95.82 56.95 72.54 74.33
Critic-RM-Rank 86.13 88.88 91.10 75.02 95.49 87.93 90.42 78.44 96.43 57.39 71.77 73.87
4.4 Evaluation on Critiques
As Critic-RM involves a crucial step for generating critiques, it is also important to evaluate the quality of critiques for target tasks. We use CriticBench to perform a comprehensive evaluation, with results detailed in Table 4. For critique accuracy, we observe that Critic-RM generates more accurate critiques compared to strong baselines, including GPT-4. This justify that Critic-RM is able to distinguish correct and flawed reasoning paths. Additionally, these critiques help the policy language model (LM) correct flawed reasoning steps, resulting in improved accuracy in refined responses. Notably, even when using the lightweight Llama-3-8B as the policy LM, the critiques guide the smaller LM to rectify initial incorrect reasoning and achieve high accuracy across five reasoning tasks.
(a) Ablation on weight scheduling.
(b) Acc. w/ different rounds.
(c) Ablation on filtering strategies.
Figure 2: Ablation studies. The y-axis is the average accuracy on RewardBench.
4.5 Ablation Studies
Effect of Two-stage Training. Figure 2(a) illustrates the performance of Critic-RM with different weight scheduling function $\lambda(t)$. The results indicate that using a constant weight across different rounds, as well as reverse weight scheduling (i.e., prioritizing reward modeling first, followed by critique generation), both negatively impact performance. Besides, Figure 2(b) shows the performance of Critic-RM with different $K$ (training epoch), where reward modeling is applied only in the final epoch. The results indicate that performance improves when $K=2$, but plateaus with further increases. Thus, $K=2$ serves as a trade-off to balance between performance and training efficiency.
Effect of Data Filtering. We further evaluate our data filtering strategy in Figure 2(c) and observe that using the entire dataset without filtering leads to poor performance, particularly in the Chat-hard domain, which requires stronger reasoning capabilities for LLMs to assess response preferences accurately. Additionally, removing noisy preference pairs improves standard reward modeling. Moreover, incorporating summarization and ranking proves to be an effective approach for boosting overall performance.
Table 5: Performance of Critic-RM and most relevant baselines with different amounts of training data.
Data Volume Method Chat Chat Hard Reasoning Safety Overall
RM 98.4 70.8 83.1 73.6 81.5
SynRM 96.9 75.1 84.2 89.1 86.3
10% CLoud 96.6 74.7 86.1 86.3 85.9
Critic-RM-summ 96.1 77.0 86.7 91.0 87.7
Critic-RM-rank 96.4 77.9 85.6 90.2 87.5
RM 98.1 69.2 85.6 81.8 83.6
SynRM 97.2 75.7 85.0 89.8 86.9
30% CLoud 97.4 76.7 86.1 87.5 86.9
Critic-RM-summ 96.9 78.7 87.4 92.2 88.8
Critic-RM-rank 97.8 77.1 86.5 93.1 88.6
RM 98.3 75.6 87.4 82.2 85.9
SynRM 97.4 76.9 85.1 90.0 87.3
50% CLoud 97.2 76.5 86.9 89.3 87.4
Critic-RM-summ 97.2 78.7 89.1 93.1 89.5
Critic-RM-rank 97.2 79.2 88.9 94.0 89.8
4.6 Data Efficiency of Reward Models
Table 5 shows the accuracy of Critic-RM and baselines on RewardBench with different volumes of training data. Overall, Critic-RM consistently outperforms the baselines across all data volumes, demonstrating superior performance even with limited labels. Notably, Critic-RM shows strong data efficiency—using just 10% of the labeled data is sufficient to surpass the standard reward model. This result highlights the data efficiency of Critic-RM, making it highly practical for real-world applications.
4.7 Case Studies
Table 6: An example that comparing Standard reward model (RM), critique-out-loud reward models (CLoud) and Critic-RM on reward prediction. Underline text denotes imprecise generation of rejected responses, Green text stands the central part that leads to accurate predictions of Critic-RM. Critic-RM can better model the preference by generating accurate critiques that identify the major strengths and weaknessses of the responses.
Prompt Describe a time when you had to make a difficult decision.
Chosen I had to make a difficult decision when I was working as a project manager at a construction company. I was in charge of a project that needed to be completed by a certain date in order to meet the client’s expectations. However, due to unexpected delays, we were not able to meet the deadline and so I had to make a difficult decision. I decided to extend the deadline, but I had to stretch the team’s resources even further and increase the budget. Although it was a risky decision, I ultimately decided to go ahead with it to ensure that the project was completed on time and that the client’s expectations were met. The project was eventually successfully completed and this was seen as a testament to my leadership and decision-making abilities.
Rejected Certainly! One situation where I had to solve a complex problem was during my time working as a project manager for a software development team. We were tasked with developing a new mobile application for a client within a tight deadline. The complex problem arose when we encountered a critical bug in the application just a few days before the scheduled release. The bug caused the app to crash randomly, making it unusable for end-users. This issue had the potential to damage our reputation and relationship with the client. To solve this problem, I immediately gathered the development team to assess the situation. We conducted a thorough investigation to identify the root cause of the bug. After extensive debugging and testing, we discovered that the issue was related to a memory leak in the code. To address the problem, we divided the team into smaller groups, each focusing on a specific aspect of the application. We implemented a rigorous testing process to identify and fix any memory leaks throughout the codebase…
RM Chosen Reward: 0.27        Rejected Reward: 0.56
CLoud Critique for Chosen Response: The assistant provides unnecessary details about the project’s context, such as the need to meet the client’s expectations, which, while relevant, could be conveyed more succinctly. Furthermore, the response focuses more on the outcome of the decision rather than the decision-making process itself. The user prompt asks for a description of a time when a difficult decision was made, implying that the user wants to understand the thought process and considerations that led to the decision. The assistant’s response glosses over these aspects, instead highlighting the success of the project and the client’s satisfaction. Additionally, the response lacks depth and introspection. Chosen Reward: 0.46 Critique for Rejected Response: The response provides a clear and specific example from the assistant’s experience, making it more relatable and engaging. The situation is well-explained, including the problem, the steps taken to address it, and the outcome. The assistant highlights their role in coordinating the team’s efforts and communicating with the client, demonstrating leadership skills. However, The response could be more concise, as some details, such as the extensive debugging and testing process, could be summarized more briefly. While the assistant mentions the importance of thorough testing and quality assurance processes, it would be beneficial to elaborate on how this was implemented in the specific situation or how it was improved upon. Rejected Reward: 0.66
Critic-RM Critique for Chosen Response: The response provides a specific scenario where a difficult decision had to be made, and it describes the decision-making process and its outcome. The response addresses the user’s input question directly and provides relevant details about the situation. The response is also clear and concise, making it easy to understand. The response does not contain any inaccuracies or inconsistencies. Chosen Reward: 0.97 Critique for Rejected Response: The response provided by the AI assistant does not accurately address the user’s input question, which asks for a time when the AI had to make a difficult decision. Instead, the response describes a situation where the AI solved a complex problem as a project manager for a software development team. The response fails to provide a personal anecdote or decision-making scenario, instead providing a hypothetical or fictional account. The response also lacks a clear decision-making process and outcome. The response would be improved by providing a more relevant and accurate answer to the user’s question. Rejected Reward: 0.07
Table 6 presents two examples from RewardBench that highlight the advantages of Critic-RM over baseline models. Standard RMs often struggle to assign higher scores to the correct response and fail to provide additional context to justify the predicted reward. While Cloud offers general feedback on response quality, it tends to focus on strengths and weaknesses without identifying the most critical aspects for evaluation. In contrast, Critic-RM accurately identifies key errors in rejected responses and explains why the chosen response is superior, leading to more accurate predictions. We list additional case studies in Appendix 8.
5 Conclusion
In this work, we introduced Critic-RM, a self-critiquing framework designed to enhance reward modeling for large language models. By harnessing LLMs’ ability to generate and refine critiques, Critic-RM implements a novel self-improvement approach that improves both critique quality and reward prediction accuracy. Experiments on multiple datasets demonstrate that Critic-RM consistently outperforms baseline reward models, showing strong data efficiency and delivering robust results even with limited labeled data. Moreover, the critiques generated by Critic-RM prove effective in helping LLMs enhance response quality. We hope that self-critiquing techniques offer a promising future direction for advancing reward modeling and improving the alignment between LLMs and human preferences.
Limitation and Future Work
Critic-RM introduces a new framework for reward modeling by leveraging self-generated critiques. While it shows promising results, several limitations exist:
Single Model Focus: Critic-RM does require the base LLM to have a certain level of critique generation ability. Testing Critic-RM across different LLM architectures could provide broader insights into its effectiveness.
Longer Inference Time: Generating critiques during inference adds computational overhead. This trade-off may affect its use in real-time applications where latency is critical for model deployment.
No Iterative Training: Critic-RM does not incorporate iterative training, where models refine themselves over multiple rounds. Adding this step could further improve reward modeling performance, as shown in recent studies (Yuan et al., 2024; Pang et al., 2024).
Acknowledgments
We would like to thank Anirudh Goyal and Thomas Scialom for the discussion on the early stage of this project.